Each of three identical bodies has a temperature-independent heat capacity C. The three bodies have initial temperatures T3>T2>T1. What is the maximum amount of work that can be extracted leaving the three bodies at a common final temperature?
1
Expert's answer
2011-11-04T12:20:42-0400


final temperature of the three bodies would be (T1+T2+T3)/3 = T1 (say)
work done in heat exchange by the three bodies = mc[(T1-T1)+(T1-T2)+(T1-T3)] = 2mc(T1+T2+T3)/3